# SUG-Occ: An Explicit Semantics and Uncertainty Guided Sparse Learning Framework for Real-Time 3D Occupancy Prediction

### v2

@@ III-A Problem Formulation
We formulate the task as predicting a voxel-wise semantic occupancy field of a 3D driving scene from a set of calibrated camera images. Given input images $I\in\mathbb{R}^{N\times H\times W\times 3}$ captured by $N$ cameras where $H$ and $W$ denote the image height and width, respectively. Each camera is associated with an intrinsic matrix $K\in SE(3)$ and an extrinsic pose $T\in SE(3)$. The objective is to infer a 3D semantic occupancy grid $Occ\in\mathbb{R}^{X\times Y\times Z}$, where $X$, $Y$, and $Z$ denote the spatial dimensions of the voxel grid along the longitudinal, lateral, and vertical axes. Each voxel in $Occ$ represents the semantic category of its corresponding spatial location, including free space as well as object categories like car, pedestrian, and vegetation.
@@ III-D Cascade Sparse Completion
Moreover, to further improve the computational efficiency, we adopt a hyper cross convolution kernel in both encoder and decoder. Unlike standard cubic convolution kernel, the hyper cross kernel aggregates features only from axial neighbors. For example, for an active voxel processed by a sparse convolution layer with kernel size $3$, convolution is performed with only $6$ axial neighbors instead of all 26 neighboring voxels. This design significantly reduces neighbor-matching operations and memory-access overhead, making sparse convolution more scalable and well-suited for real-time occupancy completion.
@@ IV-A Dataset and Metrics
We evaluate the proposed method on the widely used SemanticKITTI dataset [2], which provides dense semantic occupancy annotations by augmenting LiDAR scans from KITTI Odometry Benchmark [14], which consists of 22 outdoor driving scenarios. Following the standard setting for semantic occupancy prediction on SemanticKITTI, the prediction region is defined as a fixed 3D volume centered at the ego vehicle. The predicted region covers $51.2m$ in front of the vehicle, $25.6m$ to both left and right sides and extents from 2.0m-2.0m to $4.4m$ along the vertical axis. This region is discretized into a voxel grid with a resolution of $256\times 256\times 32$, corresponding to a voxel size of $0.2\times 0.2\times 0.2m$. Each voxel is annotated with one of $20$ classes, including 19 semantic categories and 1 free space class. In this work, we focus on monocular vision-based semantic occupancy prediction, following the setting in [42], where a single image is used as input for each frame.
@@ IV-B Implementation Details
During training, standard data augmentation strategies for depth estimation, semantic segmentation, and semantic occupancy prediction are applied. In addition, we employ a query denoising strategy with $20$ denoising queries, together with an EMA scheme for model optimization. Model is trained using the AdamW optimizer with a learning rate of $4\times 10^{-4}$ for $20$ epoches on $8$ NVIDIA H100. For efficiency evaluation, all models are measured on an NVIDIA RTX 4090 using PyTorch fp32 backend with a batch size of $1$. The FPS measurement follows the protocol in [7].

### v3

@@ III-A Problem Formulation
We formulate the task as predicting a voxel-wise semantic occupancy field of a 3D driving scene from a set of calibrated camera images. Given input images $I\in\mathbb{R}^{N\times H\times W\times 3}$ captured by NN cameras where $H$ and $W$ denote the image height and width, respectively. Each camera is associated with an intrinsic matrix $K\in SE(3)$ and an extrinsic pose $T\in SE(3)$. The objective is to infer a 3D semantic occupancy grid $Occ\in\mathbb{R}^{X\times Y\times Z}$, where $X$, $Y$, and $Z$ denote the spatial dimensions of the voxel grid along the longitudinal, lateral, and vertical axes. Each voxel in OccOcc represents the semantic category of its corresponding spatial location, including free space as well as object categories like car, pedestrian, and vegetation.
@@ III-D Cascade Sparse Completion
Moreover, to further improve the computational efficiency, we adopt a hyper cross convolution kernel in both encoder and decoder. Unlike standard cubic convolution kernel, the hyper cross kernel aggregates features only from axial neighbors. For example, for an active voxel processed by a sparse convolution layer with kernel size 33, convolution is performed with only 66 axial neighbors instead of all 26 neighboring voxels. This design significantly reduces neighbor-matching operations and memory-access overhead, making sparse convolution more scalable and well-suited for real-time occupancy completion.
@@ IV-A Dataset and Metrics
We evaluate the proposed method on the widely used SemanticKITTI dataset [2], which provides dense semantic occupancy annotations by augmenting LiDAR scans from KITTI Odometry Benchmark [14], which consists of 22 outdoor driving scenarios. Following the standard setting for semantic occupancy prediction on SemanticKITTI, the prediction region is defined as a fixed 3D volume centered at the ego vehicle. The predicted region covers $51.2m$ in front of the vehicle, $25.6m$ to both left and right sides and extents from 2.0m-2.0m to $4.4m$ along the vertical axis. This region is discretized into a voxel grid with a resolution of $256\times 256\times 32$, corresponding to a voxel size of $0.2\times 0.2\times 0.2m$. Each voxel is annotated with one of 2020 classes, including 19 semantic categories and 1 free space class. In this work, we focus on monocular vision-based semantic occupancy prediction, following the setting in [42], where a single image is used as input for each frame.
@@ IV-B Implementation Details
During training, standard data augmentation strategies for depth estimation, semantic segmentation, and semantic occupancy prediction are applied. In addition, we employ a query denoising strategy with 2020 denoising queries, together with an EMA scheme for model optimization. Model is trained using the AdamW optimizer with a learning rate of $4\times 10^{-4}$ for 2020 epoches on 88 NVIDIA H100. For efficiency evaluation, all models are measured on an NVIDIA RTX 4090 using PyTorch fp32 backend with a batch size of 11. The FPS measurement follows the protocol in [7].